The next dividend that will be paid on your stock is expected to be $2.15. Dividends grow at a constant rate of 6.5%. If the stock is currently selling for $71, calculate the expected return on this stock based on the constant growth model.
Answer #1
Expected return = (D1/P0)+g Where,
= (2.15/71.00)+0.065 D1 = $       2.15
= 9.53% P0 = $    71.00
g = 6.50%